OpenStudy (anonymous):

Find an equation of the tangent line to the curve y=sin(3x)+cos(4x) at the point (pi/6, y(pi/6)) Tangent line: y= ___ SHOW WORK! I keep getting the wrong answer.. y'= 3cos(3x)-4sin(4x) what do I do now?

OpenStudy (zepp):

Did you try to plug pi/6 in?

OpenStudy (anonymous):

ur derivative is correct.... now plug in pi/6 into your derivative to get the slope of the tangent line at x=pi/6

OpenStudy (anonymous):

yup, and I got it wrong.

OpenStudy (anonymous):

pluggin in pi/6 gives me -2sqrt(3) is that right?

OpenStudy (anonymous):

correct.... ^^^^ that's your slope

OpenStudy (anonymous):

i'll b quiet...

OpenStudy (anonymous):

I know.. and then I plug it into (y-y1)=m(x-x1)? right and plug in x and y

OpenStudy (zepp):

\[\begin{align} y'= 3\cos(3x)-4\sin(4x)\\ y'=3\cos((3)\frac{\pi}{6})-4\sin((4)(\frac{\pi}{6}))\\ y'=0-2\sqrt{3}=-2\sqrt{3} \end{align}\] The slope is correct.

OpenStudy (anonymous):

then do I plug in slope to (y-y1)=m(x-x1) or y=mx+b

OpenStudy (zepp):

\(y=mx+b\) of course!

OpenStudy (zepp):

I'll give you a hint, at which point do you want to find the equation of the tangent line? And what is this point's coordinates?

OpenStudy (anonymous):

pi/6?

OpenStudy (zepp):

A coordinate should be in this format: \((x,y)\)

OpenStudy (anonymous):

....?

OpenStudy (zepp):